# Probing false vacuum decay and bubble nucleation in a Rydberg atom array

### v1

@@ Abstract
Exploring resonant bubble nucleation. — Our focus thus far has been the the early-time exponential decay after the quench, a nucleation behavior common to classical systems, continuous quantum fields, and discrete quantum systems. However, an isolated quantum system with a discrete energy spectrum, like our Ising chain, exhibits unique long-time dynamics. First, energy conservation in the unitary evolution prevents nucleated TV bubbles from expanding indefinitely to reach the global TV. Second, the discrete spectrum gives rise to resonant conditions that can facilitate bubble nucleation.
@@
To study resonant bubble nucleation, we employ a different experimental protocol. We first prepare the Néel state, and then ramp up the Rabi frequency $\Omega$ from zero to a final value $\Omega_{f}$ over a duration $T\sim 1~\rm{\mu s}$ (see SM Sec. I). While this ramp is not strictly adiabatic, it is designed to transfer the state population into a localized region of the Hilbert space, which helps reveal the underlying nucleation resonances more clearly. Figure 4 illustrates the system’s response to this ramp protocol and the measurements of resonant bubbles. Figure 4(a) depicts the energy landscape of the product states at $V/\Delta_{l}\approx 2$, the resonance condition for $L=2$ bubbles. Each dot in the landscape represents a manifold of product states, grouped by their Hamming distance from the $|\text{Néel}\rangle$ state and their relative static energy $E/V$ (at $\Omega=0$). The color map shows the simulated projection probability of the final state, which evolves from $|\text{Néel}\rangle$ after a $T=1~\mu\text{s}$ ramp to $\Omega_{f}/2\pi=1~\text{MHz}$. The evolution, constrained by energy conservation, preferentially populates the $L=2$ bubble manifolds (located on the zero-potential-energy line). This coherence also implies that for longer evolution, the system will undergo quantum revivals, leading to oscillations in the AFM order.
@@
with site index $i\in[1,N]$ and periodic boundary conditions for the ring. As an example, Fig. 4(b) plots the measured time evolution of $\langle\hat{\Sigma}_{L}(t)\rangle$ during a ramp to $\Omega_{f}/2\pi=1.8~\text{MHz}$ under the $L=2$ resonance condition ($V/2\pi=\Delta_{g}/2\pi=10$ MHz, $\Delta_{l}/2\pi=5$ MHz). As expected, the $L=2$ bubble density shows the most significant growth, clearly outpacing the growth of $L=1$ and $L=3$ bubbles.
We define the TV bubble filling factor as $\rho=\sum_{L=1}^{N-1}(L+1)\langle\hat{\Sigma}_{L}\rangle$. As $\rho$ approaches unity, the system enters a high-bubble-density regime, where TV bubbles are separated by single-site FV background and exhibit strong interactions or blockade [PhysRevE.95.022121]. The final state in Fig. 4(b) yields an experimental (numerical) filling factor of $\rho\approx 0.46$ ($\rho\approx 0.74$), confirming that our protocol reaches this high-density limit.
This resonance is tunable: by experimentally scanning $V/\Delta_{l}$, we can selectively enhance bubbles of different lengths. Figures  4(c-e) plot the final measured bubble densities $\langle\hat{\Sigma}_{L}\rangle$ after the $\Omega$-ramp. As expected, we observe pronounced resonance peaks for $L=1$, 22, and $3$ centered near $V/\Delta_{l}\approx 1$, $2$, and $3$, respectively.
Conclusion. — In summary, we have simulated the false vacuum decay and bubble nucleation dynamics in a 1D AFM Ising ring implemented with a programmable Rydberg atom array. Exploiting single-site addressability, we studied a generalized Ising model, observing the hallmark exponential decay of the AFM order. A key finding is the strong initial-state dependence of the FVD: a proper metastable (PQG) state exhibits a stabler decay, with a rate that follows the expected exponential suppression with the inverse staggered field, whereas a simple Néel state does not. While we mainly focused on the evolution of local AFM order, future work could explore non-local observables and further clarify the role of long-range interactions in FVD dynamics. In the last part of our work, we explored the long-time dynamics, probing the resonant bubble nucleation process unique to discrete systems. Using a resonant ramp protocol, we accessed a high-bubble-density regime where bubble-bubble interactions are non-negligible. Our work pave the way for future studies of many-body tunneling on Rydberg platforms, including systems with multiple true and false vacuums that can be realized in $\mathbb{Z}_{3}$-symmetry-broken phases [Bernien_2017, Keesling_2019], as well as models of complex lattice geometries [Semeghini_2021].
Note added. — While this work is being prepared for publication, we became aware of a related effort in a 2D Rydberg array by the group of Prof. Christian Gross at the University of Tübingen [2D_Gross].
@@ Acknowledgements.
Acknowledgements. — This work is supported by the National Natural Science Foundation of China (NSFC) (Grants No. 92565306, No. 92265205, No. 12234012, and No. W2431002), the Shanghai Municipal Science and Technology Project (25LZ2601100), the Quantum Science and Technology–National Science and Technology Major Project (2021ZD0302100), and the Ministry of Science and Technology of China (Grant No. 2023YFC2205800). X. Wang at Cornell University is supported by the U.S. Department of Energy through Award Number: DE-SC0023905.
@@
Yu-Xin Chao1,2, Peiyun Ge1,2, Zhen-Xing Hua1, Chen Jia1, Xiao Wang3, Xinhui Liang1, Zongpei Yue1, Rong Lu1,
Meng Khoon Tey1,4,5, Xiao Wang2,1,6, and Li You1,2,4,5
@@
6Department of Physics, University of Oxford, UK
In this Supplemental Material, we provide additional details that support the main results of the Letter. In Section I we summarize the details on the experimental realization of staggered longitudinal field and provide the time sequences of typical experiments. In Section II we report the calibration of experimental errors, and how we include/compensate them in the numerical simulation in Fig. 3(b) in the main text. In Section III, we analyze the exponential decay behavior studied in Figs. 2 and 3 in the main text and demonstrate that the same qualitative trend is obtained using an alternative fitting procedure. In Section IV we present a BCH treatment of the AFM order and compute early-time coefficients ($\langle\hat{C}_{2}\rangle$) for Néel and PQG state, including the Transverse-Field-Ising-Model (TFIM) limit and a Hellmann–Feynman small-$\Omega$ result, alongside with the fourth-order term and higher-order structure.
@@ I.A Realization of staggered longitudinal field
In our experiment, a spatial-light-modulator (SLM) shapes an 830-nm laser beam to create an optical tweezer array for trapping the atoms. This trapping light is turned off abruptly before the main experimental sequence begins. A staggered longitudinal field, $\Delta_{\rm{stagger}}=(-1)^{j}\Delta_{l}$, is produced using a second SLM. This SLM generates a set of addressing tweezers that illuminate every other atom in the ring with a 1014-nm laser. In our setup, we address atoms on even sites (j=even), inducing a shift of $\Delta_{\rm add}=\Delta_{\uparrow}-\Delta_{\downarrow}$, where $\Delta_{\uparrow(\downarrow)}$ represents the AC Stark shifts on the $\uparrow(\downarrow)$ state. To achieve the targeted overall detuning of $\Delta_{g}+(-1)^{j}\Delta_{l}$ for the j-th atom, we set the two-photon detuning of the 1013-nm and 420-nm lasers, which are applied globally to the atoms, to $\Delta=\omega_{\rm 1013nm}+\omega_{\rm 420nm}-\omega_{\uparrow\downarrow}=-\Delta_{g}-\Delta_{l}$ and set the addressing field shift to $\Delta_{\rm add}=2\Delta_{l}$.
@@
For the first step, alignment, we adapt the protocol from Ref [thesis_Gabriel] with minor modifications. We begin by cooling the atoms near their vibrational ground state, ensuring they are localized at the center of the 830-nm tweezers. We then apply an additional blazed grating to the 1014-nm SLM’s initial phase pattern to shift the addressing tweezer positions, subsequently measuring the response of atoms. This response allows us to infer the relative position between the atoms and the 1014-nm tweezers. Our modification to the original protocol is to measure atom loss probability—by abruptly turning the addressing tweezers on and off—instead of measuring the AC Stark shift of the addressing tweezers. This loss probability exhibits a specific dependence on alignment: it is near zero for perfect overlap, increases for slight misalignment, and returns to zero when completely non-overlapped. By sweeping the blazed grating and monitoring this loss signal, we can determine the optimal addressing tweezer position. This information is fed back to generate a corrected 1014-nm SLM phase pattern. Several rounds of this feedback reduce the overall loss probability to below 3%, indicating good overlap between the two tweezer sets.
@@ I.B Experimental sequence
Whenever there is a residual offset between the center of any 1014-nm addressing tweezer and that of the 830-nm trapping tweezer, the atom, which is confined originally in the trapping tweezer, would experience a larger fluctuation in $\Delta_{\rm add}$ and also stronger heating if we abruptly turn on the addressing tweezer. To avoid this problem, we perform an “adiabatic transfer” just before the core experiment. This involves simultaneously ramping down the 830-nm tweezers while ramping up the 1014-nm addressing tweezers at the even-site until $\Delta_{\rm add}/2\pi=30$ MHz (see Fig. S1). At the end, atoms in the even sites are adiabatically transferred into the 1014-nm addressing tweezers while those in the odd sites remain at the centers of the trapping lasers.
@@ II.B Decoherence
To model the experimental system with these errors, we make use of the Lindblad master equation [BreuerPetruccione, Manzano2020]. For each site $j$ we include a decay jump operator
@@ III.A Theoretical predictions of semiclassical and quantum spin model
From the semiclassical perspective, false-vacuum decay is governed by an instanton (bounce) process [Langer1969, 1977_Coleman_Instanton, PhysRevD.16.1762]: a critical bubble of the true vacuum nucleates within the false vacuum, and the decay rate is controlled by the Euclidean action of this saddle-point trajectory. In the thin-wall limit, the action scales linearly with the critical bubble size $L_{c}\sim V/\Delta_{l}$, leading to a decay rate of the form $\Gamma\sim\frac{\Delta_{l}}{V}e^{-\lambda V/\Delta_{l}}$, with prefactor determined by fluctuations around the instanton.
References [1999_Rutkevich_PerturbativeIsingModel, Munster2000] give analytical derivations for the FV decay rate in a ferromagnetic Ising model with both transversal and global longitudinal fields. This model directly maps to our anti-ferromagnetic Ising model with both transversal and staggered longitudinal fields when $\Delta_{g}=V$. They show that in the thin-wall limit (where the critical bubble size $V/\Delta_{l}\gg 1$), the exponential decay rate of return probability (associated with creating a single bubble of resonant size) obeys
@@
The decay rate of AFM order should be considered as decay rate of return probability $\Gamma$ multiplied by the resonant bubble size $L_{c}$,
$\gamma\propto L_{c}\frac{\Delta_{l}}{V}\exp(-\lambda\frac{V}{\Delta_{l}})=\exp(-\lambda\frac{V}{\Delta_{l}}).$ (S4)
Although initially derived in the thin-wall limit, this result’s validity can be extended up to $V/\Delta_{l}\gtrsim 1.5$ by including further instanton corrections [Munster2000]. Accounting for lattice effects further refines the exponent and the prefactor predicted by the continuum theory [Lagnese2024].
@@ III.B Extraction of FV decay rate
Figure S4 shows several examples of time evolution of AFM order, from different initial states, at different staggered fields $\Delta_{l}$. Focus first on the Néel state [sub-figures (a)-(c)]. In the initial stage, the AFM order decays at a similar rate for different values of $\Delta_{l}$, with $M_{\rm{AFM}}^{\rm{~res}}(t)$ decaying from 1 to about $\rm{e}^{-0.1}$. This period is known as the quantum Zeno regime [2008_Facchi_QuantumZeno]. After the initial period, it enters into the FVD regime, which we are interested in. For Néel state however, the FVD dynamics is accompanied by persistent oscillations. This leads to difficulty in pin-pointing the FVD decay regime and in extracting the decay rate, especially when $\Delta_{l}$ is small.
@@
In Fig. 3(b) of the main text, we show that the fitted decay rate $\gamma$ of PQG state follows an excellent agreement with the theoretical prediction for a 1D ferromagnetic Ising model [Eq. (6) of ref. [2021_Lagnese_FVDinSpinChains]] for $V/\Delta_{l}\in[1,25]$. In fact, when $\Delta_{l}$ is very close to 0, the extracted $\gamma$ of the PQG state using the aforementioned fitting method deviates from that prediction. This is because the exceptionally slow and small FV decay at small $\Delta_{l}$ makes it challenging to isolate an exponential fitting window free from the post-quench oscillations [Lagnese2024]. As a result, even though the intrinsic FVD decay rate is exponentially suppressed with $V/\Delta_{l}$, these oscillations effectively inflate the fitted $\gamma$.
The mixing of FVD signal and post-quench oscillations can be mitigated by using a larger transverse field $\Omega$. A larger $\Omega$ (provided it remains sufficiently far from the critical point $2\Omega/V=1$) accelerates the FVD process, thereby allows for better extraction of the FVD signal [2021_Lagnese_FVDinSpinChains]. To show the influence of $\Omega$, we simulate the FVD of an infinite Rydberg chain with nearest-neighbor interactions for different Rabi frequencies $\Omega$ using infinite-MPS time evolution [Vidal2007iTEBD, tenpy] with bond dimension $D=100$. The results are shown in Fig. S5. For the PQG state, the $\ln\gamma$ varies linearly with respect to $V/\Delta_{l}$ for $V/\Delta_{l}\in[1,21]$ for $\Omega/V=0.3$ (corresponding to our experimental parameter). Increasing the Rabi frequency to $\Omega/V=0.35$ extends the linear regime to about $[1,35]$, whereas decreasing it to $\Omega/V=0.25$ shrinks the linear regime to about $[1,13]$. For the Néel state, all three values of $\Omega$ yield nearly the same decay behavior.
In the end, we comment that for small $\Delta_{l}$ in a finite system, the resonant bubble size may become comparable to or greater than the system’s length. Under that condition, a bubble-size blockade effect effectively forbids the nucleation [ge2025twostate].

### v2

@@ Abstract
Resonant bubble nucleation. — The early-time exponential-decay nucleation behavior we discuss so far is common to classical systems, continuous quantum fields, and discrete quantum systems. However, an isolated quantum system with a discrete energy spectrum, like our Ising chain, exhibits unique long-time dynamics. First, energy conservation in the unitary evolution prevents nucleated TV bubbles from expanding indefinitely to reach the global TV. Second, the discrete spectrum gives rise to resonant conditions that can facilitate bubble nucleation. This mechanism is akin to the resonant melting of prethermal states recently observed in a 2D Rydberg array [59], driven by coupling to specific excitation manifolds.
@@
We employ a different experimental protocol to observe the resonant nucleation. We first prepare the Néel state, and then ramp up $\Omega$ to a final value $\Omega_{f}$ over a duration $T\sim 1~\rm{\mu s}$ (see SM Sec. I [42]). While this ramp is not strictly adiabatic, it is designed to transfer the state population into a more localized region in the Hilbert space, which helps reveal the underlying nucleation resonances more clearly. Figure 4 illustrates the system’s response to this ramp protocol. Figure 4(a) depicts the energy landscape of the product states at $V/\Delta_{l}\approx 2$, the resonance condition for $L=2$ bubbles. Each dot in the landscape represents a manifold of product states, grouped by their Hamming distance from the $|\text{Néel}\rangle$ state and their relative static energy $E/V$ (at $\Omega=0$). The color map shows the simulated projection probability of the final state, which evolves from $|\text{Néel}\rangle$ after a $T=1~\mu\text{s}$ ramp to $\Omega_{f}/2\pi=1~\text{MHz}$. The evolution, constrained by energy conservation, preferentially populates the $L=2$ bubble manifolds (located on the zero-potential-energy line). This constraint implies that for longer evolution, the dynamics is dominated by the nucleation and annihilation of $L=2$ bubbles.
@@
with site index $i\in[1,N]$ and periodic boundary conditions for the ring. As an example, Fig. 4(b) plots the measured time evolution of $\langle\hat{\Sigma}_{L}(t)\rangle$ during a ramp to $\Omega_{f}/2\pi=1.8~\text{MHz}$ under the $L=2$ resonance condition ($V/2\pi=\Delta_{g}/2\pi=10$ MHz, $\Delta_{l}/2\pi=5$ MHz). As expected, the $L=2$ bubble density shows the most significant growth, clearly outpacing the growth of $L=1$ and $L=3$ bubbles. By changing $\Omega_{f}$ and $T$, we can selectively enhance bubbles of different lengths. Figures  4(c-e) plot the measured bubble densities $\langle\hat{\Sigma}_{L}\rangle$ after the $\Omega$-ramp. As expected, we observe pronounced resonance peaks for $L=1$, 22, and $3$ centered near $V/\Delta_{l}\approx 1$, $2$, and $3$, respectively.
Summary and outlook. — We simulate the false vacuum decay and bubble nucleation dynamics in a generalized 1D AFM Ising ring using a Rydberg atom array. A key finding is the strong initial-state dependence of the FVD: a proper metastable (PQG) state exhibits a cleaner and more extended exponential decay of the AFM order, with a rate that follows the expected exponential suppression with the inverse staggered field, whereas a simple Néel state does not. We also probe the resonant bubble nucleation process unique to discrete systems. In future, one may explore non-local observables and further clarify the role of long-range interactions in FVD dynamics. Our work paves the way for future studies of many-body tunneling on Rydberg platforms, including systems with multiple true and false vacuums that can be realized in $\mathbb{Z}_{3}$-symmetry-broken phases [60, 61], as well as models of complex lattice geometries [62, 63].
Note added. — During the preparation of this manuscript, we became aware of a related effort in a 2D Rydberg array by the group of Prof. Christian Gross at the University of Tübingen [64].
@@ Acknowledgements.
Acknowledgements. — This work is supported by the National Natural Science Foundation of China (NSFC) (Grants No. 92565306, No. 92265205, No. 12234012, and No. W2431002), and the Quantum Science and Technology–National Science and Technology Major Project (2021ZD0302104). R.L. is supported by the “Gravitational Wave Detection” program (2023YFC2205800) funded by the Ministry of Science and Technology of the People’s Republic of China. X.W. at Cornell University is supported by the U.S. Department of Energy through Award Number: DE-SC0023905.
@@
Yu-Xin Chao1,2,∗, Peiyun Ge1,2,∗, Zhen-Xing Hua1, Chen Jia1, Xiao Wang3, Xinhui Liang1, Zongpei Yue1, Rong Lu1,
Meng Khoon Tey1,4,5, Xiao Wang1,2,6, and Li You1,2,4,5
@@
6Department of Physics, University of Oxford, Oxfordshire, UK
In this Supplemental Material, we provide technical details supporting the experimental and theoretical results presented in the main text. Section I details the experimental implementation of the staggered longitudinal field using spatial light modulators and outlines the time sequences for state preparation and measurement. In Section II, we characterize experimental imperfections, including state preparation and measurement (SPAM) errors and decoherence, and describe their incorporation into our numerical modeling. Section III elaborates on the analysis of exponential decay dynamics, verifying the robustness of the extracted decay rates with two different fitting protocols. Finally, Section IV presents a Baker-Campbell-Hausdorff (BCH) expansion of the antiferromagnetic order parameter, providing an analytical derivation of the early-time dynamics that distinguishes the stability of the pre-quench ground (PQG) state from the Néel state.
@@ I.A Realization of staggered longitudinal field
In our setup, a spatial-light-modulator (SLM) shapes an 830-nm laser beam to create an optical tweezer array for trapping the atoms. This trapping light is turned off abruptly prior to the main experimental sequence. A staggered longitudinal field, $\Delta_{\rm{stagger}}=(-1)^{j}\Delta_{l}$, is produced using a second SLM, which creates an addressing array of 1014-nm tweezers targeting even-numbered atoms (j=even) in the ring. Due to the large detuning from the Rydberg transition ($|6P_{3/2}\rangle\rightarrow|70S_{1/2}\rangle$), the resulting AC Stark shifts of addressing tweezers, $\Delta_{\text{add}}=\Delta_{\uparrow}-\Delta_{\downarrow}$, is dominated by the ground-state component $\Delta_{\downarrow}$. For a beam waist of $1.09~\mu\text{m}$, an addressing power of $\sim 15.5$ mW yields a shift of $\Delta_{\text{add}}=30$ MHz. To achieve the targeted overall detuning of $\Delta_{g}+(-1)^{j}\Delta_{l}$ for the j-th atom, we set the two-photon detuning of the 1013-nm and 420-nm lasers, which are applied globally to the atoms, to $\Delta=\omega_{\rm 1013nm}+\omega_{\rm 420nm}-\omega_{\uparrow\downarrow}=-\Delta_{g}-\Delta_{l}$ and set the addressing shift to $\Delta_{\rm add}=2\Delta_{l}$.
@@
For the first step, alignment, we adapt the protocol from Ref [43] with minor modifications. We begin by cooling the atoms near their vibrational ground state, ensuring they are localized at the center of the 830-nm tweezers. We then apply an additional blazed grating to the 1014-nm SLM’s initial phase pattern to shift the addressing tweezer positions, subsequently measuring the response of atoms. This response allows us to infer the relative position between the atoms and the 1014-nm tweezers. Our modification to the original protocol is to measure atom loss probability—by abruptly turning the addressing tweezers on and off—instead of measuring the AC Stark shift of the addressing tweezers. This loss probability exhibits a specific dependence on alignment: it is near zero for perfect overlap, increases for slight misalignment, and returns to zero when completely non-overlapped. By sweeping the blazed grating and monitoring this loss signal, we can determine the optimal addressing tweezer position. This information is fed back to generate a corrected 1014-nm SLM phase pattern. Several rounds of this feedback reduce the overall loss probability to below 3%, indicating good overlap between the two tweezer sets.
@@ I.B Experimental sequence
Prior to the main sequence, atoms are arranged in a ring configuration with interatomic distances of $6.65$ or $7.24~\mu\text{m}$, corresponding to nearest-neighbor interactions of $10$ or $6~\text{MHz}$, respectively. The atoms are cooled to $\sim 3~\mu\text{K}$ and initialized in the ground state $|5S_{1/2},F=2,m_{F}=2\rangle$. A $30~\text{G}$ magnetic field is applied parallel to the propagation direction of the Rydberg excitation beams. Whenever there is a residual offset between the center of any 1014-nm addressing tweezer and that of the 830-nm trapping tweezer, the atom, which is confined originally in the trapping tweezer, would experience a larger fluctuation in $\Delta_{\rm add}$ and also stronger heating if we abruptly turn on the addressing tweezer. To avoid this problem, we perform an “adiabatic transfer” just before the core experiment. This involves simultaneously ramping down the 830-nm tweezers while ramping up the 1014-nm addressing tweezers at the even-site until $\Delta_{\rm add}/2\pi=30$ MHz (see Fig. S1). At the end, atoms in the even sites are adiabatically transferred into the 1014-nm addressing tweezers while those in the odd sites remain at the centers of the trapping lasers.
@@ II.B Decoherence
To model the experimental system with these errors, we make use of the Lindblad master equation [44, 45]. For each site $j$ we include a decay jump operator
@@ III.A Theoretical predictions of semiclassical and quantum spin model
From the semiclassical perspective, false-vacuum decay is governed by an instanton (bounce) process [15, 6, 7]: a bubble of the true vacuum nucleates within the false vacuum, and the decay rate is controlled by the Euclidean action of this saddle-point trajectory. In the thin-wall limit, the action scales linearly with the resonant bubble size $L_{r}\sim V/\Delta_{l}$, leading to a decay rate of the form $\Gamma\sim\frac{\Delta_{l}}{V}e^{-\lambda V/\Delta_{l}}$, with prefactor determined by fluctuations around the instanton.
References [57] give analytical derivations for the FV decay rate in a ferromagnetic Ising model with both transversal and global longitudinal fields. This model directly maps to our anti-ferromagnetic Ising model with both transversal and staggered longitudinal fields when $\Delta_{g}=V$. They show that in the thin-wall limit (where the resonant bubble size $V/\Delta_{l}\gg 1$), the exponential decay rate of return probability (associated with creating a single bubble of resonant size) obeys
@@
The decay rate of AFM order should be considered as decay rate of return probability $\Gamma$ multiplied by the resonant bubble size $L_{r}$,
$\gamma\propto L_{r}\frac{\Delta_{l}}{V}\exp(-\lambda\frac{V}{\Delta_{l}})=\exp(-\lambda\frac{V}{\Delta_{l}}).$ (S4)
Although initially derived in the thin-wall limit, this result’s validity can be extended up to $V/\Delta_{l}\gtrsim 1.5$ by including further instanton corrections [46]. Accounting for lattice effects further refines the exponent and the prefactor predicted by the continuum theory [47].
@@ III.B Extraction of FV decay rate
Figure S4 shows several examples of time evolution of AFM order, from different initial states, at different staggered fields $\Delta_{l}$. Focus first on the Néel state [sub-figures (a)-(c)]. In the initial stage, the AFM order decays at a similar rate for different values of $\Delta_{l}$, with $M_{\rm{AFM}}^{\rm{~res}}(t)$ decaying from 1 to about $\rm{e}^{-0.1}$. This period is known as the quantum Zeno regime [48]. After the initial period, it enters into the FVD regime, which we are interested in. For Néel state however, the FVD dynamics is accompanied by persistent oscillations. This leads to difficulty in pin-pointing the FVD decay regime and in extracting the decay rate, especially when $\Delta_{l}$ is small.
@@
In Fig. 3(b) of the main text, we show that the fitted decay rate $\gamma$ of PQG state follows an excellent agreement with the theoretical prediction for a 1D ferromagnetic Ising model [Eq. (6) of ref. [12]] for $V/\Delta_{l}\in[1,25]$. In fact, when $\Delta_{l}$ is very close to 0, the extracted $\gamma$ of the PQG state using the aforementioned fitting method deviates from that prediction. This is because the exceptionally slow and small FV decay at small $\Delta_{l}$ makes it challenging to isolate an exponential fitting window free from the post-quench oscillations [47]. As a result, even though the intrinsic FVD decay rate is exponentially suppressed with $V/\Delta_{l}$, these oscillations effectively inflate the fitted $\gamma$.
The mixing of FVD signal and post-quench oscillations can be mitigated by using a larger transverse field $\Omega$. A larger $\Omega$ (provided it remains sufficiently far from the critical point $2\Omega/V=1$) accelerates the FVD process, thereby allows for better extraction of the FVD signal [12]. To show the influence of $\Omega$, we simulate the FVD of an infinite Rydberg chain with nearest-neighbor interactions for different Rabi frequencies $\Omega$ using infinite-MPS time evolution [49, 50] with bond dimension $D=100$. The results are shown in Fig. S5. For the PQG state, the $\ln\gamma$ varies linearly with respect to $V/\Delta_{l}$ for $V/\Delta_{l}\in[1,21]$ for $\Omega/V=0.3$ (corresponding to our experimental parameter). Increasing the Rabi frequency to $\Omega/V=0.35$ extends the linear regime to about $[1,35]$, whereas decreasing it to $\Omega/V=0.25$ shrinks the linear regime to about $[1,13]$. For the Néel state, all three values of $\Omega$ yield nearly the same decay behavior.
In the end, we comment that for small $\Delta_{l}$ in a finite system, the resonant bubble size may become comparable to or greater than the system’s length. Under that condition, a bubble-size blockade effect effectively forbids the nucleation [51].